OpenStudy (anonymous):

What is i^486? (complex numbers)

OpenStudy (ikram002p):

humm lets try to solve ^^

OpenStudy (ikram002p):

i=sqrt -1

OpenStudy (ikram002p):

i^2=-1 ok ?

OpenStudy (amoodarya):

hint : i^4=1 i^486= i^(4(121)+2)

OpenStudy (ikram002p):

@kristina95 ur here ? :P

OpenStudy (anonymous):

so i^486 = (i^484)*(i^2) = 1*-1=-1 right?

OpenStudy (ikram002p):

i=sqrt -1 i^2=-1 486=2*243 i^2*243=i^2^243=(-1)^243 but 243 odd so i^486=-1 :D

OpenStudy (ikram002p):

yess ur right :D